# On the multiplier spectrum of polynomials

### v1

@@ 5.1. Multiplier spectrum over an arithmetic progression
We can ask to what extent the converse of Theorem 4.6 holds.
@@
Let $P(z),Q(z)\in\mathbb{C}[z]$ be (non-exceptional) polynomials of the same degree $M\geq 2$. Under what conditions do there exist positive integers $c_{1}<d$ such that
@@
for every $N\in\mathbb{Z}_{\geq 0}$? Are there examples that do not arise from Theorem 4.6, up to equivalence?
Note that in Theorem 4.6 and Question 5.1, the first term $c_{1}$ is strictly less than the common difference $d$ of the arithmetic progression. When $c_{1}\geq d$, we can always find $k\in\mathbb{Z}_{>0}$ such that $c_{1}<dk$ and then consider the arithmetic progression $(c_{1}+Ndk)_{N=0}^{\infty}$ instead. It is interesting to consider when we can take $c_{1}=d$, which relates to the following notion of stable multiplier spectrum:
@@
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. W say $f$ and $g$ have the same stable multiplier spectrum if $\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})$ for some $k\in\mathbb{Z}_{>0}$.
@@
Clearly, $f$ and $g$ have the same stable multiplier spectrum if and only if there exists an arithmetic progression $A=(Nk)_{N=1}^{\infty}$ with the first term and common difference both equal to some integer $k\in\mathbb{Z}_{>0}$ such that $S_{n}(f)=S_{n}(g)$ for every $n\in A$. Such an arithmetic progression $A=(Nk)_{N=1}^{\infty}\subseteq\mathbb{Z}_{>0}$ is called equidistant.
@@
If there exists $k\in\mathbb{Z}_{>0}$ such that $f^{\circ k}=g^{\circ k}$, then $f$ and $g$ have the same stable multiplier spectrum. For example, for $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $n\in\mathbb{Z}_{>0}$, the polynomials $P(z)=zV(z^{2n})$ and $Q(z)=-zV(z^{2n})$ (which are a specific case of Theorem 1.9 $(2)$) have the same stable multiplier spectrum, since $P^{\circ 2}=Q^{\circ 2}$.
@@
$E_{d}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C}):\exists k\in\mathbb{Z}_{>0},\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}.$
Then $E_{d}=\cup_{k\geq 1}E_{d,k}$ is a countable union of Zariski closed sets, where
$E_{d,k}=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C}):\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}$
is Zariski closed. (Note that the map
$\kappa:\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathcal{M}_{d^{k}}(\mathbb{C})\times\mathcal{M}_{d^{k}}(\mathbb{C}),([f],[g])\mapsto([f^{\circ k}],[g^{\circ k}])$
is an algebraic morphism and $E_{d,k}$ is the inverse image of $R_{d^{k},m_{d^{k}}}$ (see §1.1) under $\kappa$.) Note that for positive integers $k_{1}$ and $k_{2}$ with $k_{1}\mid k_{2}$, we have $E_{d,k_{1}}\subseteq E_{d,k_{2}}$. Hence $E_{d}$ can be written as a countable union of increasing Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})$. Currently, it is unclear whether $E_{d}$ is a Zariski closed subset.
@@
The notion of stable multiplier spectrum and the Question 5.5 are proposed by Prof. Xie and we learn them from private communications with Prof. Xie.
Replacing “multiplier” by “length” in the above discussion yields the analogous notion of stable length spectrum.
We show that the multiplier spectrum morphism over an equidistant arithmetic progression remains generically injective on the moduli space.
@@
is generically injective. Similarly, the restriction $\tilde{\tau}_{d}[n]$ of $\tau_{d}[n]$ to ${\rm MPoly}^{d}(\mathbb{C})$ is also generically injective.
@@
Let $\kappa_{d,n}:\mathcal{M}_{d}\to\mathcal{M}_{d^{n}}$ be the composition morphism $f]\mapsto[f^{\circ n}]$. Then $\tau_{d}[n]=\tau_{d^{n}}\circ\kappa_{d,n}$. By Theorem 1.3, $\tau_{d^{n}}$ is generically injective. Thus, it suffices to show $\kappa_{d,n}$ is also generically injective. This follows from [Pak25b, Theorem 5.4].
To show the generic injectivity of $\tilde{\tau}_{d}[n]$, replacing Theorem 1.3 by Theorem 1.4, it suffices to show that the composition morphism
$\tilde{\kappa}_{d,n}:{\rm MPoly}^{d}(\mathbb{C})\to{\rm MPoly}^{d^{n}}(\mathbb{C}),[f]\mapsto[f^{\circ n}]$
is also generically injective. If $d=2$, then Theorem 3.25 (2) implies the generic injectivity of $\tilde{\kappa}_{2,n}$. Assume now $d\geq 3$. By Lemmas 3.15 and 3.16, there exists a nonempty Zariski open subset $U_{d}\subseteq{\rm MPoly}^{d}(\mathbb{C})$ such that for every $\in U_{d}, ff$ is pre-simple and $G(f)=1$. We claim that $\tilde{\kappa}_{d,n}|_{U_{d}}$ is injective. Let $f],[g]\in U_{d}$ such that $\tilde{\kappa}_{d,n}([f])=\tilde{\kappa}_{d,n}([g])$. After replacing $g$ by a suitable affine conjugate if necessary, we may assume $f^{\circ n}=g^{\circ n}$. Since $f$ is pre-simple of degree $d\geq 3$, Corollary 3.9 implies that $\tau\circ f=g=f\circ\sigma$ for some $\sigma,\tau\in\mathbb{C}[z]$ of degree one. Since $G(f)=1$, we obtain $\sigma=1$, hence $f=g$. Thus, the morphism $\tilde{\kappa}_{d,n}$ is injective on $U_{d}$. Therefore, $\tilde{\tau}_{d}[n]$ is generically injective. ∎
@@
Let $d\geq 4$ be an integer. The stable multiplier map on $\mathcal{M}_{d}(\mathbb{C})$ is very generically injective; that is, there exists a subset $V\subseteq\mathcal{M}_{d}(\mathbb{C})$ of the form
@@
where each $Z_{m}$ is a proper Zariski closed subset of $\mathcal{M}_{d}(\mathbb{C})$, such that for all $f],[g]\in V$, if $f$ and $g$ have the same stable multiplier spectrum, then $[f]=[g]$.
We may ask whether the multiplier spectrum over any infinite arithmetic progression is generically injective on the moduli space for further study. (Note that generalized Lattès maps are sparse in the moduli space, and hence so are maps arising from the Ritt move process.)
@@ 5.2. Non-injective locus and length spectrum
We aim to provide a more precise description of the non-injective locus than that given in Theorem 1.9, in light of Conjecture 1.3.
@@
  • Can we bound $N$ and #S\#S in Theorem 1.9 explicitly?
@@
  • Can we describe ${\rm PNI}_{d}$ more precisely (possibly ignoring zero-dimensional components)?
  • What can be said for rational maps (not just polynomials) with the same multiplier spectrum?
There are less known results for the length spectrum than the multiplier spectrum, since the length spectrum contains less information and the failure of being an (algebraic) morphism between schemes for the length spectrum map. It is more difficult to study the length spectrum. A significant result proved by Ji and Xie proved in this direction is that aside from the flexible Lattès family, the length spectrum determines the conjugacy class of rational maps up to finitely many choices [JX23, Theorem 1.5]. We can ask some further questions for the length spectrum as follows:
  • Does there exist a Zariski-dense open subset $U$ of $\mathcal{M}_{d}(\mathbb{C})$ such that every pair $(x,y)\in U\times U$ with $L(x)=L(y)$ must have the form $([f],[f])$ or $([f],[\overline{f}])$? Here $\overline{f}$ represents the rational map obtained by applying complex conjugate to all the coefficients of $f$. Note that a positive answer to the above question is equivalent to the correctness of a conjecture of Ji-Xie [JX25, Conjecture 1.9]. We can ask the same question with the weaker requirement that $U$ has full (or just positive) Lebesgue measure (which may not be Zariski-dense open). Similar questions can be asked for polynomials as well.

### v2

@@ 5.1. Multiplier spectrum over an arithmetic progression
We could ask to what extent the converse of Theorem 4.6 holds.
@@
Let $P(z),Q(z)\in\mathbb{C}[z]$ be (non-exceptional) polynomials of the same degree $m\geq 2$. Under what conditions do there exist integers $d>c_{1}>0$ such that
@@
for all $N\in\mathbb{Z}_{\geq 0}$? Are there examples that do not arise from Theorem 4.6, up to equivalence?
Note that in Theorem 4.6 and Question 5.1, the first term $c_{1}$ of the arithmetic progression is strictly less than the common difference $d$. If $c_{1}\geq d$, then we can take $k\in\mathbb{Z}_{>0}$ with $c_{1}<dk$ and consider the arithmetic progression $(c_{1}+Ndk)_{N=0}^{\infty}$ instead.
It is interesting to consider the case $c_{1}=d$, which relates to the notion of a stable multiplier spectrum:
@@
Let $f,g\in\mathbb{C}(z)$ be two rational maps of the same degree $d\geq 2$. We say that $f$ and $g$ have the same stable multiplier spectrum and they are stably isospectral if $\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})$ for some $k\in\mathbb{Z}_{>0}$.
@@
Clearly, $f$ and $g$ are stably isospectral if and only if there exists an arithmetic progression $A=(Nk)_{N=1}^{\infty}$ with the first term and common difference equal to an integer $k>0$ such that $S_{n}(f)=S_{n}(g)$ for all $n\in A$. Such an arithmetic progression $A=(Nk)_{N=1}^{\infty}\subseteq\mathbb{Z}_{>0}$ is called equidistant.
@@
If there exists $k\in\mathbb{Z}_{>0}$ such that $f^{\circ k}=g^{\circ k}$, then $f$ and $g$ are stably isospectral. For example, for $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$ and $n\in\mathbb{Z}_{>0}$, the polynomials $P(z)=zV(z^{2n})$ and $Q(z)=-zV(z^{2n})$ are stably isospectral, since $P^{\circ 2}=Q^{\circ 2}$.
@@
$E_{d}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\colon\exists k\in\mathbb{Z}_{>0},\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}.$
Write $E_{d}=\cup_{k\geq 1}E_{d,k}$ as a countable union, where
$E_{d,k}:=\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\colon\tau_{d^{k}}(f^{\circ k})=\tau_{d^{k}}(g^{\circ k})\}.$
Note that $E_{d,k}$ is the inverse image of $R_{d^{k},m_{d^{k}}}$ (see §1.1) under the morphism
$\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})\to\mathcal{M}_{d^{k}}(\mathbb{C})\times\mathcal{M}_{d^{k}}(\mathbb{C}),([f],[g])\mapsto([f^{\circ k}],[g^{\circ k}]);$
so $E_{d,k}$ is Zariski closed. For positive integers $k_{1}$ and $k_{2}$ with $k_{1}\mid k_{2}$, we have $E_{d,k_{1}}\subseteq E_{d,k_{2}}$. Hence $E_{d}=\cup_{n\geq 1}E_{d,n!}$ is a countable union of increasing Zariski closed subsets of $\mathcal{M}_{d}(\mathbb{C})\times\mathcal{M}_{d}(\mathbb{C})$. Currently, it is unclear whether $E_{d}$ is Zariski closed.
@@
The notion of stable multiplier spectrum and Question 5.5 were proposed by Prof. Xie. We learned about these ideas from Prof. Xie through private communications.
Replacing “multiplier” with “length” in the above discussion yields the analogous notion of a stable length spectrum.
We show that the multiplier spectrum morphism over an equidistant arithmetic progression remains generically injective on the moduli space:
@@
is generically injective. Similarly, the restriction $\tilde{\tau}_{d}[n]$ of $\tau_{d}[n]$ to $\mathrm{MPoly}^{d}(\mathbb{C})$ is also generically injective.
@@
Let $\kappa_{d,n}:\mathcal{M}_{d}\to\mathcal{M}_{d^{n}}$ be the composite morphism $f]\mapsto[f^{\circ n}]$. Then $\tau_{d}[n]=\tau_{d^{n}}\circ\kappa_{d,n}$. By Theorem 1.3, $\tau_{d^{n}}$ is generically injective. Thus, it suffices to show that $\kappa_{d,n}$ is also generically injective, which follows directly from [Pak25b, Theorem 5.4].
To show the generic injectivity of $\tilde{\tau}_{d}[n]$, replacing Theorem 1.3 with Theorem 1.4 in the above argument, it suffices to show that the composite morphism
$\tilde{\kappa}_{d,n}:\mathrm{MPoly}^{d}\to\mathrm{MPoly}^{d^{n}},[f]\mapsto[f^{\circ n}]$
is generically injective. If $d=2$, then Theorem 3.31 2 implies the generic injectivity of $\tilde{\kappa}_{2,n}$. Assume now $d\geq 3$. By Lemmas 3.21 and 3.22, there exists a non-empty Zariski open subset $U_{d}\subseteq\mathrm{MPoly}^{d}(\mathbb{C})$ such that for every $\in U_{d}, ff$ is pre-simple and $G(f)=1$. We claim that $\tilde{\kappa}_{d,n}\mid_{U_{d}}$ is injective. Let $f],[g]\in U_{d}$ such that $\tilde{\kappa}_{d,n}([f])=\tilde{\kappa}_{d,n}([g])$, i.e., $f^{\circ n}]=[g^{\circ n}]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$. After conjugating $g$ by a suitable degree one polynomial if necessary, we may assume $f^{\circ n}=g^{\circ n}$. Since $g$ is pre-simple, Corollary 3.12 implies that $\tau\circ f=g=f\circ\sigma$ for some $\sigma,\tau\in\mathrm{Aff}(\mathbb{C})$. Then $\sigma=1$ because $G(f)=1$, hence $f=g$. Thus, the morphism $\tilde{\kappa}_{d,n}$ is injective on $U_{d}$. Therefore, $\tilde{\tau}_{d}[n]$ is generically injective. ∎
@@
Let $d\geq 2$ be an integer. Then the stable multiplier map on $\mathcal{M}_{d}(\mathbb{C})$ is very generically injective in the sense that there exists a subset $V\subseteq\mathcal{M}_{d}(\mathbb{C})$ of the form
@@
where each $Z_{m}$ is a proper Zariski closed subset of $\mathcal{M}_{d}(\mathbb{C})$, such that for all $f],[g]\in V$, if $f$ and $g$ are stably isospectral, then $[f]=[g]$. An analogous result holds for polynomials as well.
Note that generalized Lattès maps are sparse in the moduli space, and hence so are maps arising from Ritt moves. For further study, we may ask whether the multiplier spectrum over any infinite arithmetic progression is generically injective on the moduli space.
@@ 5.2. Non-injective locus and length spectrum
We aim to obtain a more precise description of the non-injective locus of $\tilde{\tau}_{d}$ than that given in Theorem 1.10, in light of Conjecture 1.8.
@@
  • Can we bound $N$ and #S\#S in Theorem 1.10 (more) effectively and explicitly?
@@
  • Can we describe $\mathrm{PNI}_{d}$ more precisely (possibly ignoring zero-dimensional components)?
  • What can be said for isospectral rational maps (not just polynomials)?
There are fewer known results for the length spectrum compared to the multiplier spectrum. The length spectrum is more difficult to study because it contains less information and fails to be an (algebraic) morphism between schemes. A significant result [JX23, Theorem 1.5] proved by Ji and Xie in this direction is that apart from the flexible Lattès family, the length spectrum determines the conjugacy class of rational maps up to only finitely many choices. We propose the following questions regarding the length spectrum:
  • Let $d\geq 2$ be an integer. Does there exist a non-empty Zariski open subset $U$ of $\mathcal{M}_{d}(\mathbb{C})$ such that every pair $(x,y)\in U\times U$ with $L(x)=L(y)$ must be of the forms $([f],[f])$ or $([f],[\overline{f}])$? Here $\overline{f}$ represents the rational map obtained by applying complex conjugation to all the coefficients of $f$. Note that a conjecture of Ji and Xie [JX25, Conjecture 1.9] implies that this question has a positive answer. We could ask the same question with the weaker requirement that $U$ has full (or just positive) Lebesgue measure (which may not be Zariski open). Similar questions can be asked for polynomials as well.